If (3 over 4 - 2 over 3) + (1 over 2 + 1 over 3) is calculated and the answer reduced to simplest terms, what is the denominator of the resulting fraction?
Answer by perfectdose(7)

First thing is to find the common denominator.
The LCD is 12
Multiple everything by 12/1
When you reduce with the denominator, the 'leftovers' go to the numerator. (multiple with whats there)
9/12-8/12+6/12+4/12= (9-8+6+4=11)
11/12